OpenStudy (anonymous):

All of the following are ways to write 25 percent of N except A. 0.25 N B. 25N/100 C. 1/4 N D. 25 N The answer is D but I want to know why

OpenStudy (anonymous):

In this case, N is just a variable. Let's make N 100. A. 0.25*100=25 25 is 25% of 100 so this checks out. B. 25(100)/100=25 25 is 25% of 100 so this checks out too. C. 1/4= 0.25(100)=25 25 is 25% of 100 so this checks out as well. D. 25(100)= 2500 2500 is greater than 100 so it's more than 100% of 100. As to why that is, it's because you're multiplying the variable by 25, but that won't get you 25% of the variable.